# Quantifying Statistical Significance in Diffusion-Based Anomaly Localization via Selective Inference

### v4

@@ Abstract
Anomaly localization in images—identifying regions that deviate from expected patterns—is vital in applications such as medical diagnosis and industrial inspection. A recent trend is the use of image generation models in anomaly localization, where these models generate normal-looking counterparts of anomalous images, thereby allowing flexible and adaptive anomaly localization. However, these methods inherit the uncertainty and bias implicitly embedded in the employed generative model, raising concerns about the reliability. To address this, we propose a statistical framework based on selective inference to quantify the significance of detected anomalous regions. Our method provides $p$-values to assess the false positive detection rates, providing a principled measure of reliability. As a proof of concept, we consider anomaly localization using a diffusion model and its applications to medical diagnoses and industrial inspections. The results indicate that the proposed method effectively controls the risk of false positive detection, supporting its use in high-stakes decision-making tasks.
@@ 1 Introduction
Anomaly localization using image generation models, particularly diffusion models, has shown great promise across diverse domains such as medical diagnosis and industrial inspection Li et al. (2023); Iqbal et al. (2023); Lu et al. (2023); Zhang et al. (2023); Fontanella et al. (2024); Tebbe and Tayyub (2024); Sheng et al. (2024). These models reconstruct a normal-looking version of an input image, and differences between the input and the reconstruction highlight potential anomalies. Compared to traditional methods, generative approaches are highly suitable for settings where annotations for anomalous regions are unavailable. Moreover, generative approaches can flexibly handle heterogeneity by adapting to individual images–e.g., patient-specific characteristics in medical diagnosis and product-specific traits in industrial inspection. Among various image generation models, diffusion models, in particular, offer high fidelity and stability, outperforming other methods in image quality and anomaly localization.
@@
While generative approaches offer powerful and flexible capabilities for anomaly localization, a major concern is that the inherent uncertainty and bias in generative models can affect localization performance Fithian et al. (2014); Taylor and Tibshirani (2015); Lee et al. (2016); Duy and Takeuchi (2022); Miwa et al. (2023); Shiraishi et al. (2024). These models are trained on specific datasets composed of normal images, and the quality of the generated normal-looking images depends heavily on the distribution of the dataset and how well the model has learned the underlying distribution. As a result, uncertainties or biases in the dataset or training process can cause incorrect reconstructions, leading to inaccurate localizations and misidentification of anomalies. Such risks are especially critical in high-stakes domains such as medical diagnosis and industrial inspection, where even minor errors can have serious consequences. Therefore, it is essential to incorporate rigorous uncertainty quantification framework and statistical safeguards to ensure reliable deployment in critical applications.
@@
To address this issue, we propose a statistical testing framework based on Selective Inference (SI) to assess the statistical significance of detected anomalies. SI has recently emerged as a promising approach for conducting statistical inference on hypotheses that are selected based on observed data. In this framework, inference is performed using the conditional distribution given the selection event, thereby accounting for the uncertainty and bias associated with the hypothesis selection. Following this SI framework, our key idea is to apply statistical testing to detected anomalies, conditioned on the fact that the anomalous regions were identified using a specific generative model. This allows us to quantify the statistical significance of detected anomalies as a valid $p$-value, providing a rigorous estimate of the false positive rate and offering a principled metric for reliability.
@@
In this paper, as a proof of concept for the proposed statistical testing framework, we focus on a standard denoising diffusion probabilistic model (DDPM) Ho et al. (2020); Song et al. (2022) among various diffusion-based anomaly detection (AD) methods , and its applications to medical diagnostics and industrial inspections. However, the proposed framework is readily generalizable to a broader range of diffusion model architectures and is applicable to semi-supervised AD problems in other domains.
@@
Diffusion models have been effectively utilized in anomaly localization problems Pinaya et al. (2022); Fontanella et al. (2024); Wyatt et al. (2022); Mousakhan et al. (2023). In this context, the DDPM is commonly used Ho et al. (2020); Song et al. (2022). During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise. In the test phase, the model attempts to reconstruct a new test image. If the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomalous regions. Other types of generative AI has also been used for anomalous region detection task Baur et al. (2021); Chen and Konukoglu (2018); Chow et al. (2020); Jana et al. (2022).
SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm Lee and Taylor (2014); Lee et al. (2016); Tibshirani et al. (2016), and then extended to more complex feature selection methods Yang et al. (2016); Suzumura et al. (2017); Hyun et al. (2018); Rügamer and Greven (2020); Das et al. (2021). Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks Chen and Bien (2020); Tsukurimichi et al. (2021); Tanizaki et al. (2020); Duy et al. (2022); Le Duy et al. (2024); Lee et al. (2015); Gao et al. (2022); Duy et al. (2020); Jewell et al. (2022). The fundamental idea of SI is to perform an inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data. To conduct SI, it is necessary to derive the sampling distribution of test statistic conditional on the hypothesis selection event. To the best of our knowledge, SI was applied to statistical inferences on several deep learning models Duy et al. (2022); Miwa et al. (2023); Shiraishi et al. (2024); Miwa et al. (2024), but none of them works on image generation by diffusion models.
Our contributions
The main contributions of our study are summarized as follows.  [^1] We note that our contribution is not the development of a new diffusion-based anomaly localization algorithm, but rather the introduction of a rigorous statistical testing framework designed to quantify the statistical reliability of anomalous regions identified by diffusion-based AD methods.
  • We propose a novel statistical testing framework to assess the significance of anomaly localization results derived from diffusion model-based methods, offering a principled basis for evaluating the reliability of detected anomalies.
@@
The implementation code for reproducing all experimental results is provided at https://github.com/tkatsuoka/DAL-test
Figure 1: Schematic illustration of anomaly localization on a brain MRI image dataset using a diffusion model and the proposed DAL-Test. When a test image—potentially containing an anomalous region—is fed into a trained diffusion model, its normal-looking version is generated through the forward and reverse processes. By comparing the input image with the generated normal-looking version, the anomalous region can be identified. We propose the Diffusion-based Anomaly Localization (DAL) Test, which leverages the selective inference framework to compute valid $p$-values that quantify the statistical significance of anomalous regions detected by a diffusion model, based on a test statistic defined over the input and reference images.
@@ 2 Anomaly localization by diffusion models
This section describes the anomaly localization task based on a diffusion model, which is explored as a proof of concept in this study. The process of anomaly localization using generative models can generally be divided into two phases. First, during the training phase, a denoising diffusion probabilistic model (DDPM) is trained using a dataset composed exclusively of normal images. The model learns the distribution of normal images through two key processes: the diffusion process, in which noise is gradually added to an image, and the reverse diffusion process, in which the original image is reconstructed from noise. Through this procedure, the model enhances its capacity to reconstruct normal image structures by acquiring denoising capabilities at each step. Next, during the testing phase, the reverse diffusion process is conditionally applied to an unseen input test image. In this step, the model reconstructs an image that closely resembles the input but conforms to its learned notion of “normality”, causing anomalous regions to be poorly reproduced. An anomaly score is then computed based on the difference between the reconstructed image and the input test image, and the spatial distribution of this score is analyzed to localize anomalies. By applying thresholding to the score map, anomalous regions can be clearly identified.
In this study, for the purpose of proof of concept, we adopt standard denoising diffusion models as our choice of diffusion model Ho et al. (2020); Song et al. (2022). The following outlines the image reconstruction process of a trained DDPM. Given a test image which possibly contain anomalous regions, a denoising diffusion model is used to generate the corresponding normal image. The reconstruction process consists of two processes called forward process (or diffusion process) and reverse process. In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},I)$. Let $\bm{x}$ be an image represented as a vector with each element corresponding to a pixel value. Given an original test image $\bm{x}_{0}$, noisy images $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps. We consider the distribution of the original and noisy test images, which is denoted by $q(\bm{x})$, and approximate the distribution by a parametric model $p_{\theta}(\bm{x})$ with $\theta$ being the parameters. Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2},<\dots<\beta_{T}<1$, the forward process is written as
@@
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. In the reverse process, a parametric model in the form of $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\bm{x}_{t-1};\mu_{\theta}(\bm{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\bm{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$. Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$. In DDPM Ho et al. (2020), the loss function for training the noise component is simply written as $||\epsilon^{(t)}_{\theta}(\bm{x}_{t})-\epsilon_{t}||^{2}_{2}$. Based on (2), given a noisy image $\bm{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\bm{x}_{t-1}$ is obtained as
@@
In practice, the reverse process starts from $\bm{x}_{T^{\prime}}$ with $T^{\prime}<T$. Namely, we reconstruct the original input image not from the completely noisy one, but from a one which still contains individual information of the original input image. The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image. Conversely, the larger $T^{\prime}$ results in the retention of only large scale features, thereby converting more of the anomalous regions into normal regions Ho et al. (2020); Mousakhan et al. (2023).
@@
0:  Input image $\bm{x}$
1:  $\bm{x}_{T^{\prime}}\leftarrow\sqrt{\alpha_{T^{\prime}}}\bm{x}+\sqrt{1-\alpha_{T^{\prime}}}\epsilon$
2:  for $t=T^{\prime},\dots,1$ do
3:     $f^{(t)}_{\theta}(\bm{x}_{t})\leftarrow(\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t}))/\sqrt{\alpha_{t}}$
4:     $\bm{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\bm{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t})+\sigma_{t}\epsilon_{t}$
5:  end for
5:  Reconstructed image $\bm{x}_{0}$
3 Statistical test for diffusion-based anomalous localization
In this section, we formulate the statistical test for detecting anomalous regions by a trained diffusion model. As shown in Figure 1, anomalous region detection by diffusion models is conducted as follows. First, in the training phase, the diffusion model is trained only on normal images. Then, in the test phase, we feed a test image which might contain anomalous regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$. By appropriately selecting $T^{\prime}$, we can generate a normal image that retain individual characteristics of the test input image. If the image does not contain anomalous regions, the reconstructed image is expected to be similar to the original test image. On the other hand, if the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. Therefore, the anomalous regions can be detected by comparing the original test image and its reconstructed one.
@@ 3 Statistical test for diffusion-based anomalous localization
We develop a statistical test to quantify the reliability of decision-making based on images generated by diffusion models. To develop a statistical test, we interpret an image as a sum of a true signal component $\bm{\mu}\in\mathbb{R}^{n}$ and a noise component $\bm{\varepsilon}\in\mathbb{R}^{n}$. We emphasize that the noise component $\bm{\varepsilon}$ should not be confused with the noise $\epsilon$ used in the forward process. Regarding the true signal component, each pixel can have an arbitrary value without any particular assumption or constraint. On the other hand, regarding the noise component, it is assumed to follow a Gaussian distribution, and their covariance matrix is estimated using normal data different from that used for the training of the diffusion model, which is the standard setting of SI. Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector
@@
where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix. In the following, we use capital $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image vector is denoted as $\bm{x}$.
Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}\colon\mathbb{R}^{n}\ni\bm{X}\to\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$. The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error. When identifying anomalous regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise. In this study, we simply used an averaging filter. Let us denote the averaging filter as $\mathcal{F}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$. Then, the process of obtaining the (filtered) reconstruction error is written as
@@
where absolute value is taken pixel-wise. Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $\mathcal{E}_{i}(\bm{X})$ for each pixel $i\in[n]$. Specifically, we define the anomalous region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in\mathbb{R}^{+}$, i.e.,
@@
where $\bm{\mu}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the unknown true signal vector of the reference image and the $\bm{\varepsilon}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the noise component. Then, we consider the following null and alternative hypotheses:
$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i}\\ ~{}~{}~{}\text{v.s.}~{}~{}~{}\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i},\end{gathered}$ (10)
where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference images in the observed anomalous region $\mathcal{M}_{\bm{x}}$, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different. A reasonable test statistic for the statistical test in (10) is the difference in mean pixel values between the test image and the reference image in the anomalous region $\mathcal{M}_{\bm{x}}$, i.e.,
@@
where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 00 otherwise. In this case, the $p$-value called naive $p$-value, and defined as
@@
If we can identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$, we can compute a valid $p$-value that control the false positive detection rate (i.e., the type I error rate).

### v5

# Statistical Test for Diffusion-Based Anomaly Localization via Selective Inference

@@ Abstract
Anomaly localization in images—identifying regions that deviate from normal patterns—is vital in applications such as medical diagnosis and industrial inspection. A recent trend is the use of image generation models in anomaly localization, where these models generate normal-looking counterparts of anomalous images, thereby allowing flexible and adaptive anomaly localization. However, these methods inherit the uncertainty and bias implicitly embedded in the employed generative model, raising concerns about the reliability. To address this, we propose a statistical framework based on selective inference to quantify the significance of detected anomalous regions. Our method provides $p$-values to assess the false positive detection rates, providing a principled measure of reliability. As a proof of concept, we consider anomaly localization using a diffusion model and its applications to medical diagnoses and industrial inspections. The results indicate that the proposed method effectively controls the risk of false positive detection, supporting its use in high-stakes decision-making tasks.
@@ 1 Introduction
Anomaly localization using image generation models, particularly diffusion models, has shown great promise across diverse domains such as medical diagnosis and industrial inspection (Li et al., 2023; Iqbal et al., 2023; Lu et al., 2023; Zhang et al., 2023; Fontanella et al., 2024; Tebbe and Tayyub, 2024; Sheng et al., 2024). These models reconstruct a normal-looking version of an input image, and differences between the input and the reconstruction highlight potential anomalies. Compared to traditional anomaly localization methods, generative approaches are highly suitable for settings where annotations for anomalous regions are unavailable. Moreover, generative approaches can flexibly handle heterogeneity by adapting to individual images—e.g., patient-specific characteristics in medical diagnosis and product-specific traits in industrial inspection. Among various image generation models, diffusion models, in particular, offer high fidelity and stability, often outperforming other methods in image quality and anomaly localization.
@@
While generative approaches offer powerful and flexible capabilities for anomaly localization, a major concern is that the inherent uncertainty and bias in generative models can affect localization performance (Fithian et al., 2014; Taylor and Tibshirani, 2015; Lee et al., 2016; Duy and Takeuchi, 2022; Miwa et al., 2023; Shiraishi et al., 2024). These models are trained on specific datasets composed of normal images, and the quality of the generated normal-looking images depends heavily on the dataset distribution and how accurately the model has learned it. As a result, uncertainties or biases in the dataset or training process can cause incorrect reconstructions, leading to inaccurate localizations and misidentification of anomalies. Such risks are especially critical in high-stakes domains such as medical diagnosis and industrial inspection, where even minor errors can have serious consequences. Therefore, it is essential to incorporate a rigorous uncertainty quantification framework and statistical safeguards to ensure reliable deployment in critical applications.
@@
To address this issue, we propose a statistical testing framework based on Selective Inference (SI) to assess the statistical significance of detected anomalies. SI has emerged as a promising approach for conducting statistical inference on hypotheses that are selected based on observed data. In this framework, inference is performed using the conditional distribution given the selection event, thereby accounting for the uncertainty and bias associated with the hypothesis selection. Following this SI framework, our key idea is to apply statistical testing to detected anomalies, conditioned on the fact that the anomalous regions were identified using a specific generative model. This allows us to quantify the statistical significance of detected anomalies as a valid $p$-value, providing a rigorous estimate of the false positive rate and offering a principled metric for reliability.
@@
In this paper, as a proof of concept for the proposed statistical testing framework, we focus on a standard denoising diffusion probabilistic model (DDPM) (Ho et al., 2020; Song et al., 2022) among various diffusion-based anomaly detection (AD) methods, and demonstrate its applications to medical diagnostics and industrial inspections. However, the proposed framework is readily generalizable to a broader range of diffusion model architectures and is applicable to semi-supervised AD problems in other domains.
@@
Diffusion models have been effectively utilized in anomaly localization problems (Pinaya et al., 2022; Fontanella et al., 2024; Wyatt et al., 2022; Mousakhan et al., 2023). In this context, the DDPM is commonly used (Ho et al., 2020; Song et al., 2022). During the training phase, a DDPM model learns the distribution of normal medical images by iteratively adding and then removing noise. In the test phase, the model attempts to reconstruct a new test image. If the image contains anomalous regions, such as tumors, the model may struggle to accurately reconstruct these regions, as it has been trained primarily on normal regions. The discrepancies between the original and the reconstructed image are then analyzed to identify and highlight anomalous regions. Other types of generative models have also been used for the anomaly localization task (Baur et al., 2021; Chen and Konukoglu, 2018; Chow et al., 2020; Jana et al., 2022).
SI was first introduced within the context of reliability evaluation for linear model features when they were selected using a feature selection algorithm (Lee and Taylor, 2014; Lee et al., 2016; Tibshirani et al., 2016), and then extended to more complex feature selection methods (Yang et al., 2016; Suzumura et al., 2017; Hyun et al., 2018; Rügamer and Greven, 2020; Das et al., 2021). Then, SI proves valuable not only for feature selection problems but also for statistical inference across various data-driven hypotheses, including unsupervised learning tasks (Chen and Bien, 2020; Tsukurimichi et al., 2021; Tanizaki et al., 2020; Duy et al., 2022; Le Duy et al., 2024; Lee et al., 2015; Gao et al., 2022; Duy et al., 2020; Jewell et al., 2022). The fundamental idea of SI is to perform inference conditional on the hypothesis selection event, which mitigates the selection bias issue even when the hypothesis is selected and tested using the same data. To conduct SI, it is necessary to derive the sampling distribution of the test statistic conditional on the hypothesis selection event. To the best of our knowledge, SI was applied to statistical inferences on several deep learning models (Duy et al., 2022; Miwa et al., 2023; Shiraishi et al., 2024; Miwa et al., 2024), but none of these works addresses image generation by diffusion models.
Contributions
The main contributions of our study are summarized as follows [^1] We note that our contribution is not the development of a new diffusion-based anomaly localization algorithm, but rather the introduction of a rigorous statistical testing framework designed to quantify the statistical reliability of anomalous regions identified by diffusion-based AD methods..
  • We propose a novel statistical testing framework to assess the significance of anomaly localization results obtained from diffusion model-based methods, offering a principled basis for evaluating the reliability of detected anomalies.
@@
The implementation code for reproducing all experimental results is available at https://github.com/tkatsuoka/DAL-test
Figure 1: Schematic illustration of anomaly localization using a diffusion model and the proposed DAL-Test. When a test image—potentially containing an anomalous region—is fed into a trained diffusion model, its normal-looking version is generated through the forward and reverse processes. By comparing the input image with the generated normal-looking version, the anomalous region can be identified. We propose the Diffusion-based Anomaly Localization (DAL) Test, which leverages the selective inference framework to compute valid $p$-values that quantify the statistical significance of anomalous regions detected by a diffusion model, based on a test statistic defined over the input and reference images.
@@ 2 Anomaly localization by diffusion models
This section describes the anomaly localization task based on a diffusion model, which is explored as a proof of concept in this study. The process of anomaly localization using generative models can generally be divided into two phases. First, during the training phase, a denoising diffusion probabilistic model (DDPM) is trained using a dataset composed exclusively of normal images. The model learns the distribution of normal images through two key processes: the diffusion process, in which noise is gradually added to an image, and the reverse diffusion process, in which the original image is reconstructed from noise. Through this procedure, the model enhances its capacity to reconstruct normal image structures by acquiring denoising capabilities at each step. Next, during the testing phase, the reverse diffusion process is conditionally applied to an unseen input test image. In this step, the model reconstructs an image that closely resembles the input but conforms to its learned notion of “normality”, causing anomalous regions to be poorly reproduced. An anomaly score is then computed based on the difference between the reconstructed image and the input test image, and the spatial distribution of this score is analyzed to localize anomalies. By applying thresholding to the score map, anomalous regions can be identified.
In this study, for the purpose of proof of concept, we adopt standard denoising diffusion models as our choice of diffusion model (Ho et al., 2020; Song et al., 2022). We outline the image reconstruction process of a trained DDPM below. Given a test image that possibly contains anomalous regions, a denoising diffusion model is used to generate the corresponding normal image. The reconstruction consists of two processes called forward process (or diffusion process) and reverse process. In the forward process, noise is sequentially added to the test image so that it converges to a standard Gaussian distribution $\mathcal{N}(\bm{0},I)$. Let $\bm{x}$ be an image represented as a vector with each element corresponding to a pixel value. Given an original test image $\bm{x}_{0}$, noisy images $\bm{x}_{1},\bm{x}_{2},\dots,\bm{x}_{T}$ are sequentially generated, where $T$ is the number of noise addition steps. We consider the distribution of the original and noisy test images, which is denoted by $q(\bm{x})$, and approximate the distribution by a parametric model $p_{\theta}(\bm{x})$ with $\theta$ being the parameters. Using a sequence of noise scheduling parameters $0<\beta_{1}<\beta_{2}<\dots<\beta_{T}<1$, the forward process is written as
@@
where $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{s})$. In the reverse process, a parametric model in the form of $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}(\mu_{\theta}(\bm{x}_{t},t),\beta_{t}I)$ is employed, where $\mu_{\theta}(\bm{x}_{t},t)$ is obtained by using the predicted noise component $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$. Typically, a U-Net is used as the model architecture for $\epsilon^{(t)}_{\theta}(\bm{x}_{t})$. In DDPM (Ho et al., 2020), the loss function for training the noise component is simply written as $\epsilon^{(t)}_{\theta}(\bm{x}_{t})-\epsilon_{t}||^{2}_{2}$. Based on (2), given a noisy image $\bm{x}_{t}$ after $t$ steps, the reconstruction of the image in the previous step $\bm{x}_{t-1}$ is obtained as
@@
In practice, the reverse process starts from $\bm{x}_{T^{\prime}}$ with $T^{\prime}<T$. Namely, we reconstruct the normal-looking version of the input image not from a completely noisy one, but from one that still contains individual information of the original input image. The smaller $T^{\prime}$ ensures that the reconstructed image preserves fine details of the input image. Conversely, the larger $T^{\prime}$ results in the retention of only large-scale features, thereby converting more of the anomalous regions into normal regions (Ho et al., 2020; Mousakhan et al., 2023). The overall reconstruction process of the diffusion model is summarized in Algorithm 1.
@@
1:Input image $\bm{x}$
[…]
3:for $t=T^{\prime},\dots,1$ do
4:  $f^{(t)}_{\theta}(\bm{x}_{t})\leftarrow(\bm{x}_{t}-\sqrt{1-\alpha_{t}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t}))/\sqrt{\alpha_{t}}$
5:  $\bm{x}_{t-1}\leftarrow\sqrt{\alpha_{t-1}}\cdot f^{(t)}_{\theta}(\bm{x}_{t})+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\epsilon^{(t)}_{\theta}(\bm{x}_{t})+\sigma_{t}\epsilon_{t}$
6:end for
7:Reconstructed image $\bm{x}_{0}$
3 Statistical test for diffusion-based anomaly localization
In this section, we formulate the statistical test for detecting anomalous regions by a trained diffusion model. As shown in Figure 1, anomalous region detection by diffusion models is conducted as follows. First, in the training phase, the diffusion model is trained only on normal images. Then, in the test phase, we feed a test image that might contain anomalous regions into the trained diffusion model, and reconstruct it back from a noisy image $\mathbf{x}_{T^{\prime}}$ at step $T^{\prime}<T$. By appropriately selecting $T^{\prime}$, we can generate a normal-looking image that retains individual characteristics of the test input image. If the image does not contain anomalous regions, the reconstructed image is expected to be similar to the original test image. On the other hand, if the image contains anomalous regions, such as tumors, the model may struggle to reconstruct these regions, as it has been trained primarily on normal regions. Therefore, the anomalous regions can be detected by comparing the original test image and its reconstructed one.
@@ 3 Statistical test for diffusion-based anomaly localization
We develop a statistical test to quantify the reliability of decision-making based on images generated by diffusion models. To develop a statistical test, we interpret an image as a sum of a true signal component $\bm{\mu}\in\mathbb{R}^{n}$ and a noise component $\bm{\varepsilon}\in\mathbb{R}^{n}$. We emphasize that the noise component $\bm{\varepsilon}$ should not be confused with the noise $\epsilon$ used in the forward process. No structural assumption is imposed on the signal $\bm{\mu}$. On the other hand, regarding the noise component, it is assumed to follow a Gaussian distribution, and its covariance matrix is estimated using normal data different from that used for the training of the diffusion model, which is the standard setting of SI. Namely, an image with $n$ pixels can be represented as an $n$-dimensional random vector,
@@
where $\bm{\mu}\in\mathbb{R}^{n}$ is the unknown true signal vector and $\Sigma$ is the covariance matrix. In the following, we use $\bm{X}$ to emphasize that an image is considered as a random vector, while the observed image vector is denoted as $\bm{x}$.
Let us denote the reconstruction process of the trained diffusion model in Algorithm 1 as the mapping from an input image to the reconstructed image $\mathcal{D}\colon\mathbb{R}^{n}\ni\bm{X}\to\mathcal{D}(\bm{X})\in\mathbb{R}^{n}$. The difference between the input image $\bm{X}$ and the reconstructed image $\mathcal{D}(\bm{X})$ indicates the reconstruction error. When identifying anomalous regions based on reconstruction error, it is useful to apply some filter to remove the influence of pixel-wise noise. In this study, we use an averaging filter. Let us denote the averaging filter as $\mathcal{F}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$. Then, the process of obtaining the (filtered) reconstruction error is written as
@@
where the absolute value is taken pixel-wise. Anomalous regions are then obtained by applying a threshold to the filtered reconstruction error $\mathcal{E}_{i}(\bm{X})$ for each pixel $i\in[n]$. Specifically, we define the anomalous region as the set of pixels whose filtered reconstruction error is greater than a given threshold $\lambda\in\mathbb{R}^{+}$, i.e.,
@@
where $\bm{\mu}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the unknown true signal vector of the reference image, $\bm{\varepsilon}^{\mathrm{ref}}\in\mathbb{R}^{n}$ is the noise component, and $\bm{X}^{\mathrm{ref}}$ is assumed to be independent of $\bm{X}$. Then, we consider the following null and alternative hypotheses:
$\begin{gathered}\mathrm{H}_{0}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}=\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i}\\ ~~~\text{vs.}~~~\\ \mathrm{H}_{1}\colon\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu_{i}\neq\frac{1}{|\mathcal{M}_{\bm{x}}|}\sum_{i\in\mathcal{M}_{\bm{x}}}\mu^{\mathrm{ref}}_{i},\end{gathered}$ (10)
where $\mathrm{H}_{0}$ is the null hypothesis that the mean pixel values are the same between the test image and the reference image in the observed anomalous region $\mathcal{M}_{\bm{x}}$, while $\mathrm{H}_{1}$ is the alternative hypothesis that they are different. A reasonable test statistic for the statistical test in (10) is the difference in mean pixel values between the test image and the reference image in the anomalous region $\mathcal{M}_{\bm{x}}$, i.e.,
@@
where $\bm{1}^{n}_{\mathcal{C}}\in\mathbb{R}^{n}$ is an $n$-dimensional vector whose elements are $1$ if they belong to the set $\mathcal{C}$ and 0 otherwise. In this case, the $p$-value is called the naive $p$-value, and is defined as
@@
If we can identify the sampling distribution of the test statistic $T(\bm{X},\bm{X}^{\mathrm{ref}})$, we can compute a valid $p$-value that controls the false positive detection rate (i.e., the type I error rate).